# 2.5cm2.5cm2cm2cm

### v1

@@ 2.4. Intermezzo: totally mixed Nash equilibria vs singular vector tuples
Definition 2.12 indicates that an $n$-tuple $(v^{(1)},\ldots,v^{(n)})$ is a singular vector tuple of $T$ if and only if $(\mu_{1}v^{(1)},\ldots,\mu_{n}v^{(n)})$ is also a singular vector tuple of $T$ for any nonzero $\mu_{1},\ldots,\mu_{n}\in{\mathbb{R}}$. Thus, we can unambiguously define a singular vector tuple of $T$ as the point $([v^{(1)}],\ldots,[v^{(n)}])\in{\mathbb{P}}^{\boldsymbol{d}}$. The locus of singular vector tuples of $T$ is the closed subscheme in ${\mathbb{P}}^{\boldsymbol{d}}$ defined by all the $2\times 2$ minors of the $d_{i}\times 2$ matrix $[T({\boldsymbol{v}}_{-i})\mid v^{(i)}]$ for every $i\in[n]$.
In their paper [FO14], Friedland and Ottaviani showed that the scheme of singular vector tuples of the tensor can be expressed as the zero scheme of a global section of a certain vector bundle on ${\mathbb{P}}^{\boldsymbol{d}}$ of rank $D=\dim{\mathbb{P}}^{\boldsymbol{d}}$. Furthermore, they derived a formula for the number of distinct singular vector tuples of a generic tensor by using the top Chern class of the vector bundle [FO14, Theorem 1].
Contrary to the case of totally mixed Nash equilibria stated in Theorem 2.7, a generic tensor of format ${\boldsymbol{d}}=(d_{1},\ldots,d_{n})$ always has at least one singular vector tuple. Furthermore, if $d_{1}\leq\cdots\leq d_{n}$, their number is nondecreasing in $d_{n}$ and stabilizes when $d_{n}-1=\sum_{i=1}^{n-1}(d_{i}-1)$, or when ${\boldsymbol{d}}$ is a boundary format. A geometric description of this phenomenon is described in [OSV21]. It is also worth mentioning that singular vector tuples of tensors are a special case of singular vector tuples of hyperquiver representations introduced and studied in [MNS25].
To conclude this remark, we show that both finding a totally mixed Nash equilibrium of an $n$-player game $X=(X^{(1)},\ldots,X^{(n)})\in V^{\oplus n}$ and computing a singular vector tuple of a tensor $T\in V$ can be formulated as optimization problems:
@@ 3. Nash discriminants and Nash resultants
If ${\boldsymbol{d}}=(d_{1},\ldots,d_{n})$ with $d_{1}\leq\cdots\leq d_{n}$ is balanced, i.e., if it satisfies $d_{n}-1\leq\sum_{i=1}^{n-1}(d_{i}-1)$, then a game of format ${\boldsymbol{d}}$ is anticipated to have an unexpected number of totally mixed Nash equilibria when its Nash equilibrium scheme has an unexpected number of points, or equivalently if it is nonreduced of dimension 00 (so that its reduced scheme consists of less than the expected number of points) or has a positive dimensional component (so that it contains infinitely many points). We call the variety parameterizing the games whose Nash equilibrium schemes have unexpected numbers of points the Nash discriminant variety. The primary purpose of Section 3.1 is to show that the real part of the Nash discriminant variety has codimension one. A detailed study of the geometry of the Nash discriminant variety of games for a specific format will be presented in Section 3.2.
@@ 3.1. The Nash discriminant variety
$X^{(1)}=\begin{bmatrix}[cc:cc]1&2&3&2\\ 3&1&2&3\end{bmatrix}\,,\quad X^{(2)}=\begin{bmatrix}[cc:cc]3&2&3&2\\ 2&1&3&4\end{bmatrix}\,,\quad X^{(3)}=\begin{bmatrix}[cc:cc]3&5&5&1\\ 4&1&1&3\end{bmatrix}\,.$
@@
for every $[a:b]\in{\mathbb{P}}^{1}$. Therefore, the Nash equilibrium scheme ${\mathcal{Z}}_{X}$ contains a curve isomorphic to ${\mathbb{P}}^{1}$ as an irreducible component. In particular, the triples of the vectors
@@
If $X=(X^{(1)},X^{(2)})\in V^{\oplus 2}$, then its Nash equilibrium scheme ${\mathcal{Z}}_{X}$ is defined by the following two linear systems of $d-1$ equations in $d$ variables:
@@
By Theorem 2.7, a generic game $X$ admits exactly one totally mixed Nash equilibrium, and hence $X\in\Delta(d,d)$ if and only if ${\mathcal{Z}}_{X}$ has a positive dimensional component, or equivalently the $(d-1)\times(d-1)$ minors of the coefficient matrix of at least one of the linear systems in (3.2) vanish. Therefore, the Nash discriminant variety $\Delta(d,d)$ is the union of the determinantal varieties defined by the maximal minors of these coefficient matrices.
The linear entries of each of the coefficient matrices are linearly independent. Hence, after a suitable linear change of coordinates, the determinantal variety defined by the maximal minors of each coefficient matrix can be regarded as a cone over the projective determinantal variety in ${\mathbb{P}}({\mathbb{C}}^{(d-1)\times d})$ of the generic $(d-1)\times d$ matrix of rank at most $d-2$. Therefore, Proposition 3.3 follows from [ACGH85, Chapter II.5]. ∎
@@
Let $E$ be the vector bundle on ${\mathbb{P}}^{\boldsymbol{d}}$ defined in (2.4). As explained at the beginning of Section 2.3, there exists a surjective linear map from $V^{\oplus n}$ to $H^{0}({\mathbb{P}}^{\boldsymbol{d}},E)$. This implies that the Nash discriminant variety $\Delta({\boldsymbol{d}})$ is a cone over the discriminant variety of $E$
@@
Let $E$ be the vector bundle on ${\mathbb{P}}^{\boldsymbol{d}}$ introduced in (2.4). There exists a real global section $f$ of $E$ whose zero scheme is reduced of dimension 00 and consists only of real points.
@@
Given a $[{\boldsymbol{\pi}}]=([\pi^{(1)}],\ldots,[\pi^{(n)}])\in{\mathbb{P}}^{\boldsymbol{d}}$, define the subsets $A_{1},\ldots,A_{n}$ of $F$ by $(i,j)\in A_{k}$ if and only if $\ell^{(i)}_{j,k}(\pi^{(k)})=0$. Because of the genericity of the linear forms $\ell^{(i)}_{j,k}$, the point $[{\boldsymbol{\pi}}]$ is in ${\mathcal{Z}}(f)$ precisely when the following two conditions are satisfied: (1) For every $k\in[n]$, there exists a unique subset of $\{\ell_{j,k}^{(i)}\mid\text{$i\in[n]\setminus\{k\}$, $j\in[d_{i}-1]$}\}$ with $d_{k}-1$ elements that defines $[\pi^{(k)}]$, and (2) for each $(i,j)\in F$, there exists a unique integer $k\in[n]\setminus\{i\}$ such that $(i,j)\in A_{k}$. In other words, $[{\boldsymbol{\pi}}]\in{\mathcal{Z}}(f)$ if and only if the associated subsets $A_{1},\ldots,A_{n}$ are unique and form a partition of $F$ such that $|A_{i}|=d_{i}-1$ and $F_{i}\cap A_{i}=\emptyset$ for all $i\in[n]$. The restricted partition $\{A_{1},\ldots,A_{n}\}$ of $F$ (with respect to the sets $F_{i}$) is called a block derangement in [Vid17, Section 1]. As explained in [Vid17, Section 2] or in [MM97, Section 3], the number of block derangements of $F$ is known to be equal to the integer $c({\boldsymbol{d}})$ introduced in Theorem 2.7. We conclude that the zero scheme of $f$ consists of $c({\boldsymbol{d}})$ distinct points, and hence it is reduced of dimension 00. Additionally, these points are all real because so are the coefficients of the linear forms $\ell^{(i)}_{j,k}$. ∎
@@
There exists a real global section $f$ of $E$ whose zero scheme is reduced of dimension 00 and contains at least two nonreal points.
@@
$(i)$ If $k\in\{3,\ldots,n\}$ and $(i,j)\in A_{k}=B_{k}$, define $f_{j}^{(i)}\coloneqq\prod_{r\neq i}\ell_{j,r}^{(i)}$ where $\ell^{(i)}_{j,r}$ is a generic linear form in $\pi^{(r)}$ such that $\{\ell^{(i)}_{j,k}\mid(i,j)\in A_{k}\}$ defines the point $[e_{1}^{(k)}]\in{\mathbb{P}}^{d_{k}-1}$, where $\{e_{1}^{(k)},\ldots,e_{d_{k}}^{(k)}\}$ is the standard basis for $V_{k}$.
$(ii)$ If $k\in[2]$ and $(i,j)\in A_{k}\cap B_{k}$, define $f_{j}^{(i)}\coloneqq\prod_{r\neq i}\ell_{j,r}^{(i)}$ where $\ell^{(i)}_{j,r}$ is a generic linear form in $\pi^{(r)}$, with the property that $\{\ell^{(i)}_{j,k}\mid(i,j)\in A_{k}\cap B_{k}\}$ defines the line ${\mathcal{L}}_{k}\subset{\mathbb{P}}^{d_{k}-1}$ spanned by $[e_{1}^{(k)}]$ and $[e_{2}^{(k)}]$.
@@
then the discriminant of $\widetilde{S}$ takes the negative value $-4$. Using the coefficients above and picking generic values for the remaining coefficients $c_{\alpha\beta}^{(i,j)}$ in $b_{j}^{(i)}$, define $f_{j}^{(i)}\coloneqq b_{j}^{(i)}\,\prod_{r\in\{3,\ldots,n\}\setminus\{i\}}\ell_{j,r}^{(i)}$ for every $(i,j)\in(A_{1}\cap B_{2})\cup(A_{2}\cap B_{1})$, where $\ell_{j,r}^{(i)}$ is a generic linear form in $\pi^{(r)}$.
We set $f^{(i)}\coloneqq(f_{1}^{(i)},\ldots,f_{d_{i}-1}^{(i)})\in H^{0}({\mathbb{P}}^{\boldsymbol{d}},{\mathcal{O}}({\boldsymbol{1}}_{i})^{\oplus(d_{i}-1)})$ using the components $f_{j}^{(i)}$ defined in steps $(i)-(iii)$. We show that every block derangement $C=\{C_{1},\ldots,C_{n}\}$ of $F$ can be associated with a point in the zero scheme of $f\coloneqq(f^{(1)},\ldots,f^{(n)})\in H^{0}({\mathbb{P}}^{\boldsymbol{d}},E)$:
@@
The last set of the $\dim{\mathbb{P}}^{\boldsymbol{d}}-2=D-2$ linear equations of $S_{1}$ defines the subscheme ${\mathcal{L}}_{1}\times{\mathcal{L}}_{2}\times\prod_{k=3}^{n}\{[e_{1}^{(k)}]\}\subset{\mathbb{P}}^{\boldsymbol{d}}$. Restricting the first subsystem of $S_{1}$ consisting of the two bilinear equations to this subscheme yields the bilinear system $\widetilde{S}$ studied before. In particular, the solution set of $S_{1}$ is the union of two pairwise conjugate points of ${\mathbb{P}}^{\boldsymbol{d}}$.
@@
$S_{2}\colon\begin{cases}b_{j}^{(i)}=0&\text{for every $k\in[2]$ and $(i,j)\in C_{k}\cap[(A_{1}\cap B_{2})\cup(A_{2}\cap B_{1})]$}\\[2.0pt] \ell_{j,k}^{(i)}=0&\text{for every $k\in[n]$ and $(i,j)\in C_{k}\setminus[(A_{1}\cap B_{2})\cup(A_{2}\cap B_{1})]$.}\end{cases}$
The last set of the $D-2$ linear equations of $S_{2}$ defines the subscheme ${\mathcal{L}}_{1}^{\prime}\times{\mathcal{L}}_{2}^{\prime}\times\prod_{k=3}^{n}\{p^{(k)}\}\subset{\mathbb{P}}^{\boldsymbol{d}}$, where ${\mathcal{L}}_{k}^{\prime}\subset{\mathbb{P}}^{d_{k}-1}$ is the line defined by $\{\ell^{(i)}_{j,k}\mid(i,j)\in C_{k}\setminus[(A_{1}\cap B_{2})\cup(A_{2}\cap B_{1})]\}$ for all $k\in[2]$ and $p^{(k)}\in{\mathbb{P}}^{d_{k}-1}$ is the point defined by $\{\ell^{(i)}_{j,k}\mid(i,j)\in C_{k}\setminus[(A_{1}\cap B_{2})\cup(A_{2}\cap B_{1})]\}$ for all $k\in\{3,\ldots,n\}$. The intersection between ${\mathcal{L}}_{1}^{\prime}\times{\mathcal{L}}_{2}^{\prime}\times\prod_{k=3}^{n}\{p^{(k)}\}$ and the subset cut out by the first two bilinear equations of $S_{2}$ consists of two distinct points, both different from the solutions of $S_{1}$. One of the solutions of $S_{2}$ is the point of ${\mathbb{P}}^{d}$ associated with the block derangement $C$.
$(3)$ Suppose that $[(A_{1}\cap B_{2})\cup(A_{2}\cap B_{1})]\subset C_{k}$ for some $k\in[2]$. This case holds only if $d_{k}\geq 3$. We consider the case $k=1$ for the sake of simplicity. Consider the system
@@
In this case, the last set of the $D-2$ linear equations of $S_{3}$ defines the subscheme ${\mathcal{N}}\times\prod_{k=2}^{n}\{q^{(k)}\}\subset{\mathbb{P}}^{\boldsymbol{d}}$, where ${\mathcal{N}}$ is a plane in ${\mathbb{P}}^{d_{1}-1}$ and $q^{(k)}$ is a point in ${\mathbb{P}}^{d_{k}-1}$ for all $k\in\{2,\ldots,n\}$. The intersection between ${\mathcal{N}}\times\prod_{k=2}^{n}\{q^{(k)}\}$ and the subset cut out by the first two bilinear equations of $S_{3}$ is the point in ${\mathbb{P}}^{\boldsymbol{d}}$ associated with the block derangement $C$.
$(4)$ Suppose that $|C_{1}\cap[(A_{1}\cap B_{2})\cup(A_{2}\cap B_{1})]|=1$ and $C_{2}\cap[(A_{1}\cap B_{2})\cup(A_{2}\cap B_{1})]=\emptyset$. Similarly, one considers the case when the sets $C_{1}$ and $C_{2}$ are swapped. Consider the system
@@
In this case, the last set of the $D-1$ linear equations of $S_{4}$ defines the subscheme ×${\mathcal{L}}\times\prod_{k=2}^{n}\{r^{(k)}\}\subset{\mathbb{P}}^{\boldsymbol{d}}$, where ${\mathcal{L}}$ is a line in ${\mathbb{P}}^{d_{1}-1}$ and $r^{(k)}$ is a point in ${\mathbb{P}}^{d_{k}-1}$ for all $k\in\{2,\ldots,n\}$. The intersection between ×${\mathcal{L}}\times\prod_{k=2}^{n}\{r^{(k)}\}$ and the subset cut out by the first bilinear equation of $S_{4}$ is the point in ${\mathbb{P}}^{\boldsymbol{d}}$ associated with the block derangement $C$.
@@
Since $E$ is globally generated, the dependency locus ${\mathcal{C}}$ of $f_{1}$ and $f_{2}$ is a nonsingular curve [Ein82, Lemma 2.5]. We show the irreducibility of ${\mathcal{C}}$ and find the genus $g({\mathcal{C}})$ of ${\mathcal{C}}$ by utilizing the Eagon-Northcott complex of the sheaf morphism $(f_{1},f_{2})\colon{\mathcal{O}}^{\oplus 2}\to E$

### v2

@@ 2.4. Intermezzo: totally mixed Nash equilibria vs singular vector tuples
Definition 2.13 indicates that an $n$-tuple $(v^{(1)},\ldots,v^{(n)})$ is a singular vector tuple of $T$ if and only if $(\mu_{1}v^{(1)},\ldots,\mu_{n}v^{(n)})$ is also a singular vector tuple of $T$ for any nonzero $\mu_{1},\ldots,\mu_{n}\in{\mathbb{R}}$. Thus, we can unambiguously define a singular vector tuple of $T$ as the point $([v^{(1)}],\ldots,[v^{(n)}])\in{\mathbb{P}}^{\boldsymbol{d}}$. The locus of singular vector tuples of $T$ is the closed subscheme in ${\mathbb{P}}^{\boldsymbol{d}}$ defined by all the $2\times 2$ minors of the $d_{i}\times 2$ matrix $T({\boldsymbol{v}}_{-i})\mid v^{(i)}]$ for every $i\in[n]$.
In their paper [FO14], Friedland and Ottaviani showed that the scheme of singular vector tuples of the tensor can be expressed as the zero scheme of a global section of a certain vector bundle on ${\mathbb{P}}^{\boldsymbol{d}}$ of rank $D=\dim{\mathbb{P}}^{\boldsymbol{d}}$. Furthermore, they derived a formula for the number of distinct singular vector tuples of a generic tensor by using the top Chern class of the vector bundle [FO14, Theorem 1]. It is also worth mentioning that singular vector tuples of tensors are a special case of singular vector tuples of hyperquiver representations introduced and studied in [MNS25].
We show that both finding a totally mixed Nash equilibrium of an $n$-player game $X=(X^{(1)},\ldots,X^{(n)})\in V^{\oplus n}$ and computing a singular vector tuple of a tensor $T\in V$ can be formulated as optimization problems:
@@
To conclude this remark, we highlight a relevant difference between totally mixed Nash equilibria and singular vector tuples. A consequence of [FO14, Theorem 1] is that a generic tensor of format ${\boldsymbol{d}}=(d_{1},\ldots,d_{n})$ always has at least one singular vector tuple. Furthermore, if $d_{1}\leq\cdots\leq d_{n}$, their number is nondecreasing in $d_{n}$ and stabilizes when $d_{n}-1=\sum_{i=1}^{n-1}(d_{i}-1)$, or when $\boldsymbol{d}$} is a boundary format. A geometric description of this phenomenon is described in [OSV21]. Instead, Theorem 2.8 implies that a generic game of format $\boldsymbol{d}$} with $d_{n}-1>\sum_{i=1}^{n-1}(d_{i}-1)$ does not admit totally mixed Nash equilibria.
@@ 3. Nash discriminants and Nash resultants
If ${\boldsymbol{d}}=(d_{1},\ldots,d_{n})$ with $d_{1}\leq\cdots\leq d_{n}$ is balanced, i.e., if it satisfies $d_{n}-1\leq\sum_{i=1}^{n-1}(d_{i}-1)$, then a game of format $\boldsymbol{d}$} is anticipated to have an unexpected number of totally mixed Nash equilibria when its Nash equilibrium scheme has an unexpected number of points, or equivalently if it is nonreduced of dimension 0 (so that its reduced scheme consists of less than the expected number of points) or has a positive dimensional component (so that it contains infinitely many points). We call the variety parameterizing the games whose Nash equilibrium schemes have unexpected numbers of points the Nash discriminant variety. The primary purpose of Section 3.1 is to show that the real part of the Nash discriminant variety has codimension one. A detailed study of the geometry of the Nash discriminant variety of games for a specific format will be presented in Section 3.2.
@@ 3.1. The Nash discriminant variety
$X^{(1)}=\begin{bmatrix}[cc:cc]1&3&2&1\\ 3&2&2&3\end{bmatrix}\,,\quad X^{(2)}=\begin{bmatrix}[cc:cc]3&2&2&1\\ 3&3&2&4\end{bmatrix}\,,\quad X^{(3)}=\begin{bmatrix}[cc:cc]3&4&5&1\\ 5&1&1&3\end{bmatrix}\,$.
@@
for every $a:b]\in{\mathbb{P}}^{1}$. Therefore, the Nash equilibrium scheme ${\mathcal{Z}}_{X}$ contains a curve isomorphic to ${\mathbb{P}}^{1}$ as an irreducible component. In particular, the triples of the vectors
@@
If $X=(X^{(1)},X^{(2)})\in V^{\oplus 2}$, then its Nash equilibrium scheme ${\mathcal{Z}}_{X}$ is defined by the following two linear systems of d1d-1 equations in $d$ variables:
@@
By Theorem 2.8, a generic game $X$ admits exactly one totally mixed Nash equilibrium, and hence $X\in\Delta(d,d)$ if and only if ${\mathcal{Z}}_{X}$ has a positive dimensional component, or equivalently the $(d-1)\times(d-1)$ minors of the coefficient matrix of at least one of the linear systems in (3.2) vanish. Therefore, the Nash discriminant variety $\Delta(d,d)$ is the union of the determinantal varieties defined by the maximal minors of the coefficient matrices of the linear systems in (3.2)
$\begin{pmatrix}x_{11}^{(1)}-x_{21}^{(1)}&\cdots&x_{1d}^{(1)}-x_{2d}^{(1)}\\ \vdots&&\vdots\\ x_{11}^{(1)}-x_{d1}^{(1)}&\cdots&x_{1d}^{(1)}-x_{dd}^{(1)}\end{pmatrix},\,\begin{pmatrix}x_{11}^{(1)}-x_{21}^{(1)}&\cdots&x_{1d}^{(1)}-x_{2d}^{(1)}\\ \vdots&&\vdots\\ x_{11}^{(1)}-x_{d1}^{(1)}&\cdots&x_{1d}^{(1)}-x_{dd}^{(1)}\end{pmatrix}$.
Note that the linear entries of each of the coefficient matrices are linearly independent, as explained in the proof of Proposition 2.7. Hence, after a suitable linear change of coordinates, the determinantal variety defined by the maximal minors of each coefficient matrix can be regarded as a cone over the projective determinantal variety in ${\mathbb{P}}({\mathbb{C}}^{(d-1)\times d})$ of the generic $(d-1)\times d$ matrix of rank at most d2d-2. Therefore, Proposition 3.3 follows from [ACGH85, Chapter II.5]. ∎
@@
Let $E$ be the vector bundle on ${\mathbb{P}}^{\boldsymbol{d}}$ defined in (2.4). As explained in Proposition 2.7, there exists a surjective linear map from $V^{\oplus n}$ to $H^{0}({\mathbb{P}}^{\boldsymbol{d}},E)$. This implies that the Nash discriminant variety $\Delta({\boldsymbol{d}})$ is a cone over the discriminant variety of $E$
@@
Let $E$ be the vector bundle on ${\mathbb{P}}^{\boldsymbol{d}}$ introduced in (2.4). There exists a real global section $f$ of $E$ whose zero scheme is reduced of dimension 0 and consists only of real points.
@@
Given a ${\boldsymbol{\pi}}]=([\pi^{(1)}],\ldots,[\pi^{(n)}])\in{\mathbb{P}}^{\boldsymbol{d}}$, define the subsets $A_{1},\ldots,A_{n}$ of $F$ by $(i,j)\in A_{k}$ if and only if $\ell^{(i)}_{j,k}(\pi^{(k)})=0$. Because of the genericity of the linear forms $\ell^{(i)}_{j,k}$, the point $\boldsymbol{\pi}$}] is in ${\mathcal{Z}}(f)$ precisely when the following two conditions are satisfied: (1) For every $k\in[n]$, there exists a unique subset of $\{\ell_{j,k}^{(i)}\mid\text{$i\in[n]\setminus\{k\}$, $j\in[d_{i}-1]$}\}$ with $d_{k}-1$ elements that defines $\pi^{(k)}]$, and (2) for each $(i,j)\in F$, there exists a unique integer $k\in[n]\setminus\{i\}$ such that $(i,j)\in A_{k}$. In other words, ${\boldsymbol{\pi}}]\in{\mathcal{Z}}(f)$ if and only if the associated subsets $A_{1},\ldots,A_{n}$ are unique and form a partition of $F$ such that $|A_{i}|=d_{i}-1$ and $F_{i}\cap A_{i}=\emptyset$ for all $i\in[n]$. The restricted partition $\{A_{1},\ldots,A_{n}\}$ of $F$ (with respect to the sets $F_{i}$) is called a block derangement in [Vid17, Section 1]. As explained in [Vid17, Section 2] or in [MM97, Section 3], the number of block derangements of $F$ is known to be equal to the integer $c({\boldsymbol{d}})$ introduced in Theorem 2.8. We conclude that the zero scheme of $f$ consists of $c({\boldsymbol{d}})$ distinct points, and hence it is reduced of dimension 0. Additionally, these points are all real because so are the coefficients of the linear forms $\ell^{(i)}_{j,k}$. ∎
@@
There exists a real global section $f$ of $E$ whose zero scheme is reduced of dimension 0 and contains at least two nonreal points.
@@
$(i)$ If $k\in\{3,\ldots,n\}$ and $(i,j)\in A_{k}=B_{k}$, define $f_{j}^{(i)}\coloneqq\prod_{r\neq i}\ell_{j,r}^{(i)}$ where $\ell^{(i)}_{j,r}$ is a generic linear form in $\pi^{(r)}$ such that $\{\ell^{(i)}_{j,k}\mid(i,j)\in A_{k}\}$ defines the point $e_{1}^{(k)}]\in{\mathbb{P}}^{d_{k}-1}$, where $\{e_{1}^{(k)},\ldots,e_{d_{k}}^{(k)}\}$ is the standard basis for $V_{k}$.
$(ii)$ If $k\in[2]$ and $(i,j)\in A_{k}\cap B_{k}$, define $f_{j}^{(i)}\coloneqq\prod_{r\neq i}\ell_{j,r}^{(i)}$ where $\ell^{(i)}_{j,r}$ is a generic linear form in $\pi^{(r)}$, with the property that $\{\ell^{(i)}_{j,k}\mid(i,j)\in A_{k}\cap B_{k}\}$ defines the line ${\mathcal{L}}_{k}\subset{\mathbb{P}}^{d_{k}-1}$ spanned by $e_{1}^{(k)}]$ and $e_{2}^{(k)}]$.
@@
then the discriminant of $\widetilde{S}$ takes the negative value 4-4. Using the coefficients above and picking generic values for the remaining coefficients $c_{\alpha\beta}^{(i,j)}$ in $b_{j}^{(i)}$, define $f_{j}^{(i)}\coloneqq b_{j}^{(i)}\,\prod_{r\in\{3,\ldots,n\}\setminus\{i\}}\ell_{j,r}^{(i)}$ for every $(i,j)\in(A_{1}\cap B_{2})\cup(A_{2}\cap B_{1})$, where $\ell_{j,r}^{(i)}$ is a generic linear form in $\pi^{(r)}$.
We set $f^{(i)}\coloneqq(f_{1}^{(i)},\ldots,f_{d_{i}-1}^{(i)})\in H^{0}({\mathbb{P}}^{\boldsymbol{d}},{\mathcal{O}}({\boldsymbol{1}}_{i})^{\oplus(d_{i}-1)})$ using the components $f_{j}^{(i)}$ defined in steps (i)(iii)(i)-(iii). We show that every block derangement $C=\{C_{1},\ldots,C_{n}\}$ of $F$ can be associated with a point in the zero scheme of $f\coloneqq(f^{(1)},\ldots,f^{(n)})\in H^{0}({\mathbb{P}}^{\boldsymbol{d}},E)$:
@@
The last set of the $\dim{\mathbb{P}}^{\boldsymbol{d}}-2=D-2$ linear equations of $S_{1}$ defines the subscheme ${\mathcal{L}}_{1}\times{\mathcal{L}}_{2}\times\prod_{k=3}^{n}\{[e_{1}^{(k)}]\}\subset{\mathbb{P}}^{\boldsymbol{d}}$. Restricting the first subsystem of $S_{1}$, consisting of the two bilinear equations, to this subscheme yields the bilinear system $\widetilde{S}$ previously studied. In particular, the solution set of $S_{1}$ is the union of two pairwise conjugate points of ${\mathbb{P}}^{\boldsymbol{d}}$.
@@
[…]
The last set of the D2D-2 linear equations of $S_{2}$ defines the subscheme ${\mathcal{L}}_{1}^{\prime}\times{\mathcal{L}}_{2}^{\prime}\times\prod_{k=3}^{n}\{p^{(k)}\}\subset{\mathbb{P}}^{\boldsymbol{d}}$, where ${\mathcal{L}}_{k}^{\prime}\subset{\mathbb{P}}^{d_{k}-1}$ is the line defined by $\{\ell^{(i)}_{j,k}\mid(i,j)\in C_{k}\setminus[(A_{1}\cap B_{2})\cup(A_{2}\cap B_{1})]\}$ for all $k\in[2]$ and $p^{(k)}\in{\mathbb{P}}^{d_{k}-1}$ is the point defined by $\{\ell^{(i)}_{j,k}\mid(i,j)\in C_{k}\setminus[(A_{1}\cap B_{2})\cup(A_{2}\cap B_{1})]\}$ for all $k\in\{3,\ldots,n\}$. The intersection between ${\mathcal{L}}_{1}^{\prime}\times{\mathcal{L}}_{2}^{\prime}\times\prod_{k=3}^{n}\{p^{(k)}\}$ and the subset cut out by the first two bilinear equations of $S_{2}$ consists of two distinct points, both different from the solutions of $S_{1}$. One of the solutions of $S_{2}$ is the point of ${\mathbb{P}}^{d}$ associated with the block derangement $C$.
$(3)$ Suppose that $(A_{1}\cap B_{2})\cup(A_{2}\cap B_{1})]\subset C_{k}$ for some $k\in[2]$. This case holds only if $d_{k}\geq 3$. We consider the case $k=1$ for the sake of simplicity. Consider the system
@@
In this case, the last set of the D2D-2 linear equations of $S_{3}$ defines the subscheme ${\mathcal{N}}\times\prod_{k=2}^{n}\{q^{(k)}\}\subset{\mathbb{P}}^{\boldsymbol{d}}$, where $\mathcal{N}$} is a plane in ${\mathbb{P}}^{d_{1}-1}$ and $q^{(k)}$ is a point in ${\mathbb{P}}^{d_{k}-1}$ for all $k\in\{2,\ldots,n\}$. The intersection between ${\mathcal{N}}\times\prod_{k=2}^{n}\{q^{(k)}\}$ and the subset cut out by the first two bilinear equations of $S_{3}$ is the point in ${\mathbb{P}}^{\boldsymbol{d}}$ associated with the block derangement $C$.
$(4)$ Suppose that $|C_{1}\cap[(A_{1}\cap B_{2})\cup(A_{2}\cap B_{1})]|=1$ and $C_{2}\cap[(A_{1}\cap B_{2})\cup(A_{2}\cap B_{1})]=\emptyset$. Similarly, one considers the case where the sets $C_{1}$ and $C_{2}$ are swapped. Consider the system
@@
In this case, the last set of the D1D-1 linear equations of $S_{4}$ defines the subscheme ${\mathcal{L}}\times\prod_{k=2}^{n}\{r^{(k)}\}\subset{\mathbb{P}}^{\boldsymbol{d}}$, where $\mathcal{L}$} is a line in ${\mathbb{P}}^{d_{1}-1}$ and $r^{(k)}$ is a point in ${\mathbb{P}}^{d_{k}-1}$ for all $k\in\{2,\ldots,n\}$. The intersection between ${\mathcal{L}}\times\prod_{k=2}^{n}\{r^{(k)}\}$ and the subset cut out by the first bilinear equation of $S_{4}$ is the point in ${\mathbb{P}}^{\boldsymbol{d}}$ associated with the block derangement $C$.
@@
The two preceding lemmas serve as a preparation for Theorem 3.7. The main idea of its proof is as follow. Consider two distinct sections $f_{1},f_{2}$ in $H^{0}({\mathbb{P}}^{\boldsymbol{d}},E)$ with real coefficients in the complement of the real part of $\Delta(E)$, hence ${\mathcal{Z}}(f_{i})$ is a zero-dimensional reduced complex scheme of ${\mathbb{P}}^{\boldsymbol{d}}$ such that $|{\mathcal{Z}}(f_{i})\cap{\mathbb{P}}^{\boldsymbol{d}}_{\mathbb{R}}|<\infty$ for all $i\in\{1,2\}$. If $f_{1}$ and $f_{2}$ are connected by a path $\{f_{t}\mid t\in[1,2]\}\subseteq H^{0}({\mathbb{P}}^{\boldsymbol{d}},E)$ of sections with real coefficients, such that $0\in{\mathbb{R}}^{D}$ is a regular value of $f_{t}\colon{\mathbb{P}}^{\boldsymbol{d}}_{\mathbb{R}}\to{\mathbb{R}}^{D}$ for all $t\in[1,2]$, then Thom’s Isotopy Lemma ensures that the topology of ${\mathcal{Z}}(f_{t})\cap{\mathbb{P}}^{\boldsymbol{d}}_{\mathbb{R}}$ is preserved along the parameter space $[1,2]$. This implies that if additionally $|{\mathcal{Z}}(f_{1})\cap{\mathbb{P}}^{\boldsymbol{d}}_{\mathbb{R}}|=|{\mathcal{Z}}(f_{2})\cap{\mathbb{P}}^{\boldsymbol{d}}_{\mathbb{R}}|$, then $|{\mathcal{Z}}(f_{t})\cap{\mathbb{P}}^{\boldsymbol{d}}_{\mathbb{R}}|=|{\mathcal{Z}}(f_{1})\cap{\mathbb{P}}^{\boldsymbol{d}}_{\mathbb{R}}|$ for all $t\in[1,2]$. Therefore, the path $\{f_{t}\mid t\in[1,2]\}$ does not meet the real part of $\Delta(E)$. In the upcoming proof, we instead start with two distinct sections $f_{1},f_{2}$ in $H^{0}({\mathbb{P}}^{\boldsymbol{d}},E)$ such that $|{\mathcal{Z}}(f_{1})\cap{\mathbb{P}}^{\boldsymbol{d}}_{\mathbb{R}}|\neq|{\mathcal{Z}}(f_{2})\cap{\mathbb{P}}^{\boldsymbol{d}}_{\mathbb{R}}|$. We apply Thom’s Isotopy Lemma to show that every path $\{f_{t}\mid t\in[1,2]\}$ must intersect the real part of $\Delta(E)$. In particular, the complement of the latter subset is not path-connected, which implies that the real part of $\Delta(E)$ has codimension one.
@@
Since $E$ is globally generated as explained at the beginning of Section 2.3, the dependency locus $\mathcal{C}$} of $f_{1}$ and $f_{2}$ is a nonsingular curve [Ein82, Lemma 2.5]. We show the irreducibility of $\mathcal{C}$} and find the genus $g({\mathcal{C}})$ of $\mathcal{C}$} by utilizing the Eagon-Northcott complex of the sheaf morphism $(f_{1},f_{2})\colon{\mathcal{O}}^{\oplus 2}\to E$